OpenStudy (anonymous):

tripling the greater of two consecutive even integers gives the same result as subtracting 10 from the lesser even integer. what are the integers?

OpenStudy (anonymous):

hello saifoo!

OpenStudy (saifoo.khan):

Integer A: x Integer B : x +1 3(x+1) = x - 10 Please check @satellite73

OpenStudy (saifoo.khan):

And Hello there Satellite. :)

OpenStudy (anonymous):

oops consecutive "even" integers, so if one is \(x\) the next one is \(x+2\)

OpenStudy (saifoo.khan):

Ops, my bad.

OpenStudy (anonymous):

lol close though

OpenStudy (anonymous):

\(3(x+2)=x-10\) solve for \(x\) in maybe 4 steps

OpenStudy (saifoo.khan):

Correction: Integer A: x Integer B : x +2 3(x+2) = x - 10

OpenStudy (saifoo.khan):

This is my problem from the day i was born. I never read full questions. -_-